Determine the field of existence of the function f(x) = 1 / √(1-3x) 


1
Expert's answer
2021-08-31T05:04:41-0400
"\begin{matrix}
 1-3x\geq0\\
 1-3x\not=0
\end{matrix}"

"x<\dfrac{1}{3}"

Domain: "(-\infin, \dfrac{1}{3})."


Range: "(0, \infin)"